charges -q, Q and -q are placed at equal distance on a straight line, if the potential energy of the system of three charges is zero, then find ratio of Q/q?

Grade:12th pass

1 Answers

Eshan
5 years ago
Let the separation between charges -q and Q bex.
So the separation between the other pair of -q and Q is alsox, and that between -q and -q is2x.

Since the potential energy of the whole system is zero,

-\dfrac{kQq}{x}-\dfrac{kQq}{x}+\dfrac{kq^2}{2x}=0
\implies \dfrac{kq^2}{2x}=\dfrac{2kQq}{x}\implies \dfrac{Q}{q}=\dfrac{1}{4}